The license plates for a certain state consist of two letters followed by a five - digit number such that the first digit of the number is not zero. An example for two letters and four digits is PK-2446. (If the problem has no solution, enter NONE in the answer box.)
(a) How many different license plates can be produced?
(b) How many different plates do not have a repeated letter?

(c) How many plates do not have any repeated digits in the number part of the plate?

(d) How many plates do not have a repeated letter and also do not have any repeated digits?

Answer by stanbon(75887)

The license plates for a certain state consist of two letters followed by a five - digit number such that the first digit of the number is not zero. An example for two letters and four digits is PK-2446. (If the problem has no solution, enter NONE in the answer box.)
(a) How many different license plates can be produced?
# = 26*26*9*10^4 =60,840,000
---------------------------------------------------
(b) How many different plates do not have a repeated letter?
# = 26*25*9*10^4
--------------------------------------
(c) How many plates do not have any repeated digits in the number part of the plate?
# = 26^2*9*9*8*7
---------------------------
(d) How many plates do not have a repeated letter and also do not have any repeated digits?
# = 26*26*9*9*8*7
=====================
Cheers,
Stan H.